Compute the perfectly competitive equilibrium quantity if inverse demand is given by

P = 700 - 12Q

and firms face constant marginal cost of

mc = 100
Answer #1

P=700-12Q

mc= 100

IN PERFECTLY COMPETITIVE MARKET P=MC

THUS

700-12Q=100

(ADDING 12Q ON BOTH SIDES AND SUBTRACTING 100 ON BOTH SIDES WE GET)

600=12Q

Q=600/12

Q=50

P= 700-12*50

P=700-600

P=100

Quantity is equal to 50 and price is equal to 100 in perfect competition